You are receiving $52,000 at the end of 15 years at 12% per year. The present value of this amount at moment zero is then:

Group of answer choices

$5,090

$9,500

None of these

$5,900
Answer #1
Dear student thankyou for using homeworklib.
Please refer the solution below.
Statement showing computation
Present value= Future value*1/((1+rate)^number of years)
Future value= $52000
rate= 12%
number of year=15
Present value= 52000*1/((1+0.12)^15)
Present value= 52000*0.182696261= $9500
So correct answer is $9500